The active element of a certain laser is made of a glass rod 33.0 cm long and 1.30 cm in diameter. Assume the average coefficient of linear expansion of the glass is equal to 9.00 10-6 (°C)−1. The temperature of the rod increases by 70.0°C. (a) What is the increase in its length?(b) What is the increase in its diameter?
cm
(c) What is the increase in its volume?
cm3
Answer #1

L = 33cm , d =1.3 cm, r =0.65cm

\alpha = 9x 10-6 (°C)−1

\DeltaT = 70 0C

(a)\alpha =\DeltaL/L(\DeltaT)

\DeltaL = \alphaL\DeltaT = (9x10-6)(0.33)(70)

\DeltaL =0.0208 cm

(b) \alpha =\Deltad/d(\DeltaT)

\Deltad = \alphad\DeltaT = (9x10-6)(0.013)(70)

\Deltad =0.000819 cm

(c) V =pi*r2L = (3.14*0.65*0.65*33) =43.78 cm^3

Coefficient of volume expansion \gamma =3\alpha = \DeltaV/V(\DeltaT)

\DeltaV= V(3\alpha)(\DeltaT) = (43.78x3x9x10-6x70)

\DeltaV=0.08274 cm3